OpenStudy (mathmagicain):

What is the reciprical of 5 3/4

OpenStudy (mathmagicain):

Will MEDAL!!!

OpenStudy (3mar):

you mean\[5\frac{ 3 }{ 4 }\]?

OpenStudy (mathmagicain):

YES

OpenStudy (mathmagicain):

I tink its 3/17 but im nfs

OpenStudy (3mar):

"nfs"?

OpenStudy (mathmagicain):

this is my first question btw so i dont really know how to use open study

OpenStudy (mathmagicain):

oh not for sure

OpenStudy (3mar):

Welcome at Openstudy!

OpenStudy (mathmagicain):

thanckx

OpenStudy (mathmagicain):

you do k12?

OpenStudy (erib):

0.17391304347

OpenStudy (3mar):

ok how do you deal with \[a \frac{ b}{ c }\rightarrow \frac{ d }{ c }\]?

OpenStudy (mathmagicain):

oh thankx will medal!

OpenStudy (3mar):

You know, Medals don't matter. What matters is that got the idea and can solve such this yourself.

OpenStudy (mathmagicain):

how do i fan?

563blackghost (563blackghost):

@erib you can not be giving the answers away, you must help the owlet so that they may understand the question and how to solve. This is in violation of the code of conduct. Please refrain from this in the future.

OpenStudy (3mar):

@mathmagicain Can we finish the problem then anything comes? WHere did you reach?

OpenStudy (mathmagicain):

ok

OpenStudy (mathmagicain):

bye